Problem 13-1A Stockholders' equity transactions and analysis LO C2, P1

Kinkaid Co. is incorporated at the beginning of this year and engages in a number of transactions. The following journal entries impacted its stockholders’ equity during its first year of operations.

General Journal Debit Credit
a. Cash 300,000
Common Stock, $25 Par Value 250,000
Paid-In Capital in Excess of Par Value, Common Stock 50,000
b. Organization Expenses 150,000
Common Stock, $25 Par Value 125,000
Paid-In Capital in Excess of Par Value, Common Stock 25,000
c. Cash 43,000
Accounts Receivable 15,000
Building 81,500
Notes Payable 59,500
Common Stock, $25 Par Value 50,000
Paid-In Capital in Excess of Par Value, Common Stock 30,000
d. Cash 120,000
Common Stock, $25 Par Value 75,000
Paid-In Capital in Excess of Par Value, Common Stock 45,000


Required:
2. How many shares of common stock are outstanding at year-end?
3. What is the amount of minimum legal capital (based on par value) at year-end?
4. What is the total paid-in capital at year-end?
5. What is the book value per share of the common stock at year-end if total paid-in capital plus retained earnings equals $695,000?

How many shares of common stock are outstanding at year-end? What is the amount of minimum legal capital (based on par value) at year-end? What is the total paid-in capital at year-end?

2. Number of outstanding shares 20,000
3. Minimum legal capital $500,000
4. Total paid-in capital
Answer #1
2
Number of outsanding shares 20000 =(250000+125000+50000+75000)/25
3
Minimum legal capital 500000 =20000*25
4
Total paid-in Capital 650000 =500000+50000+25000+30000+45000
5
Book value per common share
Choose Numerator / Choose Denominator = Book value per common share
Stockholders' Equity applicable to Common shares / Number of common shares outstanding = Book value per common share
695000 / 20000 = 34.75